OpenStudy (anonymous):

i will give medal and become fan Find the surface area of a conical grain storage tank that has a height of 43 meters and a diameter of 14 meters. Round the answer to the nearest square meter. (1 point) 959 m2 1,113 m2 2,072 m2 2,533 m2

OpenStudy (campbell_st):

is the height of the cone, the slant height or the perpendicular height...?

OpenStudy (anonymous):

\[SA = \pi (rs) +\pi r ^{2}\]\[s = \sqrt{h ^{2}-r ^{2}}\]\[s=40.657\] SA = 1049 which is none of the answers..

OpenStudy (campbell_st):

here is the difference... and you need slant height for the surface area,